# Dominant vertices and attractors’ landscape for Boolean networks

### v2

@@ Abstract.
In previous works, we introduced the notion of dominant vertices. This is a set of nodes in the underlying network whose evolution determines the whole network’s dynamics after a transient time. In this paper, we focus on the case of Boolean networks. We define a reduced graph on the dominant vertices and an induced automata network on this graph, which we prove is asymptotically equivalent to the original Boolean dynamics. Asymptotic conjugacy ensures that the systems, restricted to their respective attractors, are dynamically equivalent. For a significant class of networks, the induced automata network is indeed a reduction of the original system. In these cases, the reduction, which is obtained from the structure of dominant vertices, supplies a more tractable system with the same structure of attractors as the original one. Furthermore, the structure of the induced system allows us to establish bounds on the number and period of the attractors, as well as on the reduction of the basin’s sizes and transient lengths. We illustrate this reduction by considering a class of networks, which we call clover networks, whose dominant set is a singleton. To get insight into the structure of the basins of attraction of Boolean networks with a single dominant vertex, we complement this work with a numerical exploration of the behavior of a parametrized ensemble of systems of this kind.
Keywords: Boolean networks, Dominant vertices, Attractors’ landscape, Asymptotic equivalence.
@@ 1. Introduction
In [19], we showed that in a dissipative dynamical system defined over directed networks, the whole dynamics is determined, after a transient time, by the projection of the orbits on a subset of nodes we called dominant vertices. As we argue below, the same effect takes place in a Boolean network, where, due to the finiteness of the local rules (which are typically many-to-one), the dynamics can in general be considered as dissipative. It is natural then to ask whether a reduced dynamical system can be defined on a network of the dominant vertices only, which would eventually contain all the relevant information of the system. In this paper, we define a reduced graph on the dominant vertices and an induced automata network on it, and we prove that this induced system is asymptotically equivalent to the original Boolean dynamics. The asymptotic equivalence, which we introduce here, relates systems that are dynamically equivalent once restricted to their respective attractors. Under some conditions on the underlying network, the induced automata network is indeed a reduction of the original system. In this situation, the induced automata network is a more tractable version of the original Boolean network that preserves the structure of attractors. Furthermore, the nature of the induction allows us to establish bounds on the number and period of the attractors, as well as on the reduction of the basin’s sizes and transient lengths, depending only on the structure of the dominant set. To illustrate all this notions and explore the relation between structure of dominant vertices and the landscape of attractors, we consider a class of Boolean networks with a single dominant vertex, which we call clover networks. We complement this work with a numerical exploration of the behavior of a parametrized ensemble of systems of this kind.
@@
The exponential growth of the configuration space in Boolean networks has motivated the search for more simplified versions of the original network while retaining all its important dynamical features. This has led to several model reduction proposals, from which we single out [21, 27, 24], where some transitional nodes are suppressed once their effect is taken into account by precomputing their possible output, and [29, 28], where sets of nodes defining stable motifs are assumed to attain a constant value and then replaced by their influence on the rest of the network. It is this kind of forward preprocessing that we use when defining the induced automata network in Section 3. There are other unrelated reduction techniques, which consider other particularities of the dynamics, as for instance [4], where the reduction is achieved by identification of backward equivalent nodes.
@@ 2. Preliminaries
Consider a directed graph with vertices in the finite set $V$, arrows in the set $A\subset V\times V$, and for each $v\in V$, let $I(v):=\{u\in V:(u,v)\in A\}$ be the input set for $v$. We extend this notation to sets of nodes as follows: $I(U):=\bigcup_{u\in U}I(u)$ for each $U\subset V$. For each $s\in{\mathbb{N}}_{0}$, define recursively $I^{s+1}(U)=I(I^{s}(U))$, with $I^{0}(U)\equiv U$. A Boolean network is a finite dynamical system defined on such a directed graph. To ensure the well-definedness of the dynamics, we will always assume that the input set of every vertex $v\in V$ is non-empty.
Consider the set $\mathtt{B}:=\{-1,1\}\subset\mathbb{Z}$ and for each $x\in\mathtt{B}^{V}$ and $U\subset V$, denote by $x_{U}\in\mathtt{B}^{U}$ the restriction of the configuration $x$ to the vertices in $U$. In the case of a singleton, we will use the simpler notation $x_{u}$ instead of $x_{\{u\}}$. A Boolean network is a dynamical system on $\mathtt{B}^{V}$, generated by iteration of a function $F:\mathtt{B}^{V}\to\mathtt{B}^{V}$, with the following structure. For each $v\in V$, there exists a local rule $\phi_{v}:\mathtt{B}^{I(v)}\to\mathtt{B}$, such that $F(x)_{v}=\phi_{v}(x_{I(v)})$. We will refer to the graph $(V,A)$ used in the definition of $F$, as the underlying network. In order to simplify the notation, given $U\subset V$, we will use $\phi_{U}(x_{I(U)})$ to denote the configuration $\left(\phi_{u}\left(x_{I(u)}\right)\right)_{u\in U}\in\mathtt{B}^{U}$.
@@
The couple $(\mathtt{B}^{V},F)$ is a finite dynamical system whose behavior can be encompassed in a directed graph with vertices in $\mathtt{B}^{V}$ and arrows $x\mapsto y$ for each couple $(x,y)\in\mathtt{B}^{V}\times\mathtt{B}^{V}$ such that $F(x)=y$. This directed graph $\mathcal{T}_{F}$ is the transition diagram of the dynamical system. It is easy to verify that each connected component $C$ of $\mathcal{T}_{F}$ consists of a single cycle decorated with a certain number of directed trees rooted at a point in the cycle. Each cycle codifies a periodic attractor of $(\mathtt{B}^{V},F)$, and the connected component containing this cycle codifies its entire basin of attraction. In this context, the attractors’ landscape of $(\mathtt{B}^{V},F)$ is nothing but the structure of its transition diagram $\mathcal{T}_{F}$. We will characterize this structure by a list containing, for each connected component $\mathcal{C}$, the length of the cycle $P(\mathcal{C})$ (i.e., the period of the attractor), the size of the component $|\mathcal{C}|$ (i.e., the size of the basin of attraction), the mean length $\langle\tau\rangle_{\mathcal{C}}$ and the maximal length $\tau_{\max}(\mathcal{C})$ of the attached trees.
@@
Consider a network $(V,A)$ such that $I(v)\neq\emptyset$ for each $v\in V$. We say that $U\subset V$ determines the set $U^{\prime}\subset V$, if $I(U^{\prime})\subset U$, i.e., if the input set of each node in $U^{\prime}$ is completely contained in $U$. Given $U\subset V$, with $\partial U$ we denote the maximal set (in the sense of contention) determined by $U$, i.e., $\partial U=U^{\prime}\Leftrightarrow(I(U^{\prime})\subset U\,\wedge\,I(W)\subset U\Rightarrow W\subset U)$. The set $U\subset V$ is dominant if the chain $U_{0}:=U\subset U_{1}\subset\cdots\subset U_{d}:=V$, where $U_{\ell+1}=U_{\ell}\cup\partial U_{\ell}$ for each $0\leq\ell\leq d$. We call $U_{0}\subsetneq\cdots\subsetneq U_{d}$ the chain determined by $U$. Given a dominant set $U$, its chain is uniquely determined and we refer to its length $d=d(U)$, as the depth of the dominant set $U$. Alternatively, we can characterize the dominant sets is as follows.
@@
Figure 1. A five-vertex network with a dominant set of two vertices colored in blue, and depth $d=2$. The reduced graph, connecting only the dominant vertices and defined below, is shown aside.
@@ 3. The induced system
    Since the dominant set is not unique, the induce network could in principle depend on the choice of the dominant set. Natural choices of dominant sets are dominant sets of minimal cardinality, but even in that case, two dominant sets with the same minimal cardinality could have different recurrence lengths and non-equivalent reduced networks.
@@
    The eventual conjugacy is an equivalent relation between dynamical systems on finite sets. Reflexivity and transitivity are obvious. To prove symmetry, consider the following:
@@
Consider the Boolean network $(\mathtt{B}^{V},F)$ with underlying graph $(V,A)$. Let $U\subset V$ be a dominant set with depth $d=d(U)$, and recurrence length $\ell=\ell(U)$. Let $\left(\mathtt{B}_{\ell}^{U},\mathcal{F}\right)$ be the induced automata network, and $\mathtt{h}:\mathtt{B}^{V}\to\mathtt{B}_{\ell}^{U}$ the eventual conjugacy between both systems. Recall that ${\rm Per}_{P}(F)$ denotes the set of all $F$-periodic points of the period $P$ and ${\rm Per}(F)=\bigcup_{P\in\mathbb{N}}{\rm Per}_{P}(F)$. For each $x\in(\mathtt{B}^{V},F)$, with $\tau_{F}(x)$ we denote the transient time of the orbit starting at $x$, that is,
@@
As mentioned above, the landscape of the attractors comprises a finite number of basins attached to the periodic attractors, each corresponding to a connected component of the transition diagram $\mathcal{T}_{F}$. For each $x\in\mathtt{B}^{V}$, denote by $\mathcal{C}_{F}(x)$ the basin of attraction containing $x$. Finally, let us denote by $N_{\mathcal{A}}$ the number of attractors of $(\mathtt{B}^{V},F)$, or equivalently, the number of connected components of $\mathcal{T}_{F}$. The same notions and corresponding notations apply to the induced automata network $(\mathtt{B}_{\ell}^{U},\mathcal{F})$. We have the following results.
@@ 4. Clover networks with signed majority rule
A local rule of this kind is a signed majority rule, with signs given by the arrow map $\sigma$. In this case, to each directed cycle $C:=v_{0}\mapsto v_{1}\mapsto\cdots\mapsto v_{m}\mapsto v_{0}$ in $(V,A)$, we associate the sign
@@
Let $(V,A)$ be a clover network and $\sigma:A\to\mathtt{B}$ the interaction signs. Let $\mathcal{S}:=(\sum_{|C|=t}\sigma(C):1\leq t\leq\ell)$ be the vector formed by its cycles’ sign sums, grouped according to their length. Then, the corresponding Boolean network with signed majority rule induces an automata network such that $\mathcal{F}\left(y^{0}y^{1}\cdots y^{\ell-1}\right)=\left(\Phi(y)y^{0}y^{1}\cdots y^{\ell-1}\right)$, with $\Phi:\mathtt{B}_{\ell}\to\mathtt{B}$ such that
@@
The recursion stops at $t=\ell-1$, with $\ell=\max\{|C|:\,C\text{ is a cycle}\}$. The clover structure ensures that each $u\in I(v_{0})$ appears only in one cycle $C_{u}:=v_{0}\mapsto\cdots\mapsto u\mapsto v_{0}$, and for that cycle, the recursion above gives $x_{u}^{0}=\sigma(C_{u})y^{|C_{u}|-1}$, from which the proposition follows. $\Box$
@@
    The dynamics $F:\mathtt{B}^{V}\to\mathtt{B}^{V}$, defined by Equation (7), commutes with the symmetry $x\leftrightarrow-x$ in $\mathtt{B}^{V}$. This derives from the fact that ${\rm Sign}(-x|\,-y)=-{\rm Sign}(x,y)$ for each $x\in\mathbb{Z},\,y\in\mathtt{B}$. The induced automata, $\mathcal{F}:\mathtt{B}_{\ell}\to\mathtt{B}_{\ell}$ inherits this symmetry.
@@
    The symmetry $x\leftrightarrow-x$ reflects in a symmetry in the transition diagram for both, the original and the induced automata network.
@@
    Suppose that $\mathcal{S}$ is such that $|\mathcal{S}_{\tau}|>\sum_{t\neq\tau}^{\ell}|\mathcal{S}_{t}|$ for some $1\leq\tau\leq\ell$. In this case the induced automata network is such that $y\mapsto\left({\rm Sign}(\mathcal{S}_{\tau}\,y^{\tau-1})y^{0}\cdots y^{\ell-2}\right)$ and we have two possibilities:
@@
In what follows we analyse the landscape of attractors in an ensemble of clover Boolean networks with signed majority rule. We compare the mean structure of the attractor’s landscape of the original Boolean network to that of induced automata network. The aim is to evaluate the influence of the sign distribution and the distribution of cycle lengths on key dynamical features such as the number of attractors, their periods, the size of their respective basins, and the transient behavior preceding convergence. At the same time we will study the effect of network reduction, on the transient behavior. In this way we can asses who loose are the theoretical bounds established in Corollary 1, concerning the transient lengths and the size of the basins of attraction.
@@
In this way we obtain a clover network with cycles $1\mapsto n_{k}\mapsto(n_{k}+1)\mapsto\cdots\mapsto(n_{k+1}-1)\mapsto 1$, for $1\leq k\leq s+1$. The parameter $p$ controls the number and the size of the of cycles of the clover network, and we will refer to it as the folding probability. Indeed, the number $N_{c}$ of cycles is distributed according to $\mathbb{P}(N_{c}=k)=1+f(k,N-2,p)$, with f(k,N2,p)f(k,N-2,p) the binomial distribution with N2N-2 trials and parameter $p$. Hence, the expected number of cycles is $\mathbb{E}(N_{c})=1+p(N-2)$. The length $\ell_{1}$ of the cycle $C_{1}$ is such that $\mathbb{P}(\ell_{1}=\ell)=p^{\chi_{<N}(\ell)}\,(1-p)^{\ell-2}$, with $\chi_{<N}$ is the characteristic function of the set $\{\ell<N\}$. This gives an expected length $\mathbb{E}(\ell_{1})=\left(-p^{2}-(1-p)^{N}\right)/\left(p(1-p)\right)\rightarrow(1+p)/p$, as $N\to\infty$. This asymptotically coincides with the common length of the cycles in a clover network with 1+p(N2)1+p(N-2) cycles of the same length. We can estimate the distribution of the maximal length $\max_{1\leq t\leq\ell}\ell_{t}$ by considering that our random structure is reasonably modelled by a clover network with $N_{c}=1+p(N-2)$ cycles whose lengths are i.i.d. random variables with exponential distribution $\mathbb{P}(\ell_{k}=\ell)=p(1-p)^{\ell-2}$. For this simplified model,
@@
    According to Equation (11), the influence of the inhibition probability becomes less important as the size of the cycles grows. Hence, the large coordinates of $\mathcal{S}$ can be modelled by centred random variables with a variance tending to zero.
@@
    The commun number $N_{\mathcal{A}}$ of connected components of $\mathcal{T}_{F}$ and $\mathcal{T}_{\mathcal{F}}$, i.e., the commun number of attractors.
@@
    The commun length $P(\mathcal{C})$ of the cyclic core of each component of $\mathcal{T}_{F}$ and $\mathcal{T}_{\mathcal{F}}$, i.e., the commun period of the corresponding periodic attractors.
@@
    The reduction after induction of the mean depth $\Delta\tau=\langle\tau_{F}\rangle_{\mathcal{C}_{F}}-\langle\tau_{\mathcal{F}}\rangle_{\mathcal{C}_{\mathcal{F}}}$ of each component, which is noting but the difference between the basins’ mean transients before and after induction.
@@
    The reduction after induction of the maximal depth $\Delta\tau_{\max}=\max_{x\in\mathcal{C}}\tau_{F}(x)-\max_{y\in\mathtt{h}(\mathcal{C})}\tau_{\mathcal{F}}(y)$ of each component, which is noting but the difference between the basins’ maximal transients before and after induction.
We register as well, the recurrence length $\ell$, of the undelying clover network, wich serves a key parameter for the theoretical upper bounds established in Corollary 1.
We performed 500 simulations with $N=10$, for each $p,q\in\{0.3,0.6,0.9\}$, observing some interesting features regarding the attractors’ landscape.
@@
    The number $N_{\mathcal{A}}$ of attractors was, in all of our simulations, much smaller than the theoretical upper bound $\sum_{P=1}^{P_{\max}}2^{P}/P$, which indicates that the number of attractors of a given period, $N_{\mathcal{A}}(P)$, is usually smaller than the theoretical upper bound $2^{P}/P$. Besides, $N_{\mathcal{A}}$ seems to decrease with the folding probability $p$, consistent with the fact that $N_{\mathcal{A}}$ increases with the recurrence length $\ell$.
@@
    0.3 7.5240 2.8700 2.4220
    0.6 7.8800 2.7860 2.2380
    0.9 7.6440 3.0220 2.9620
    $\ell$ 5.8700 4.0660 2.6680
    Table 1. The number of attractors, $N_{\mathcal{A}}$, averaged over 500 simulations, as a function of the folding and inhibition probabilities. In the last line, the recurrence length averaged over the same 500 numerical experiments, as a function of the folding probability. In all these exploratory simulations were performed with $N=10$ nodes.

### v3

@@ Abstract.
In previous works, we introduced the notion of dominant vertices in the context of dynamical systems on networks. This is a set of nodes in the underlying network whose evolution determines the whole network’s dynamics after a transient time. In this paper, we focus on the case of Boolean networks. We define a reduced graph on the dominant vertices and an induced dynamics on this graph, which we prove is asymptotically equivalent to the original Boolean dynamics. Asymptotic conjugacy ensures that the systems, restricted to their respective attractors, are dynamically equivalent. For a significant class of networks, the induced dynamics is indeed a reduction of the original system. In these cases, the reduction, which is obtained from the structure of dominant vertices, supplies a more tractable system with the same structure of attractors as the original one. Furthermore, the structure of the induced system allows us to establish bounds on the number and period of the attractors, as well as on the reduction of the basin’s sizes and transient lengths. We illustrate this reduction by considering a class of networks, which we call clover networks, whose dominant set is a singleton. To get insight into the structure of the basins of attraction of Boolean networks with a single dominant vertex, we complement this work with a numerical exploration of the behavior of a parametrized ensemble of systems of this kind.
Dedicated to the memory of Claudine Chaouiya
@@ 1. Introduction
In [20], we showed that in a dissipative dynamical system defined over directed networks, the whole dynamics is determined, after a transient time, by the projection of the orbits on a subset of nodes we called dominant vertices. As we argue below, the same effect takes place in a Boolean network, where, due to the finiteness of the local rules (which are typically many-to-one), the dynamics can in general be considered as dissipative. It is natural then to ask whether a reduced dynamical system can be defined on a network of the dominant vertices only, which would eventually contain all the relevant information of the system. In this paper, we define a reduced graph on the dominant vertices and an induced dynamics on it, and we prove that this induced system is asymptotically equivalent to the original Boolean dynamics. The asymptotic equivalence, which we introduce here, relates systems that are dynamically equivalent once restricted to their respective attractors. Under some conditions on the underlying network, the induced network is indeed a reduction of the original system. In this situation, the induced network is a more tractable version of the original Boolean network that preserves the structure of attractors. Furthermore, the nature of the induction allows us to establish bounds on the number and period of the attractors, as well as on the reduction of the basin’s sizes and transient lengths, depending only on the structure of the dominant set. To illustrate all this notions and explore the relation between structure of dominant vertices and the landscape of attractors, we consider a class of Boolean networks with a single dominant vertex, which we call clover networks. We complement this work with a numerical exploration of the behavior of a parametrized ensemble of systems of this kind.
@@
The exponential growth of the configuration space in Boolean networks has motivated the search for more simplified versions of the original network while retaining all its important dynamical features. This has led to several model reduction proposals, from which we single out [22, 28, 25], where some transitional nodes are suppressed once their effect is taken into account by precomputing their possible output, and [30, 29], where sets of nodes defining stable motifs are assumed to attain a constant value and then replaced by their influence on the rest of the network. It is this kind of forward preprocessing that we use when defining the induced network in Section 3. There are other unrelated reduction techniques, which consider other particularities of the dynamics, as for instance [4], where the reduction is achieved by identification of backward equivalent nodes.
@@ 2. Preliminaries
Consider a directed graph with vertices in the finite set $V$, and arrows in the set $A\subset V\times V$, and for each $v\in V$, let $I(v):=\{u\in V:(u,v)\in A\}$ be the input set for $v$. We extend this notation to sets of nodes as follows: $I(U):=\bigcup_{u\in U}I(u)$ for each $U\subset V$. For each $s\in{\mathbb{N}}_{0}$, define recursively $I^{s+1}(U)=I(I^{s}(U))$, with $I^{0}(U)\equiv U$. A Boolean network is a finite dynamical system defined on such a directed graph. To ensure the well-definiteness of the dynamics, we will always assume that the input set of every vertex $v\in V$ is non-empty.
Consider the set $\mathtt{B}:=\{-1,1\}\subset\mathbb{Z}$ and for each $x\in\mathtt{B}^{V}$ and $U\subset V$, denote by $x_{U}\in\mathtt{B}^{U}$ the restriction of the configuration $x$ to the vertices in $U$. In the case of a singleton, we will use the simpler notation $x_{u}$ instead of $x_{\{u\}}$. A Boolean network is a dynamical system on $\mathtt{B}^{V}$, generated by the iteration of a function $F:\mathtt{B}^{V}\to\mathtt{B}^{V}$, with the following structure. For each $v\in V$, there exists a local rule $\phi_{v}:\mathtt{B}^{I(v)}\to\mathtt{B}$, such that $F(x)_{v}=\phi_{v}(x_{I(v)})$. We will refer to the graph $(V,A)$ used in the definition of $F$, as the underlying network. In order to simplify the notation, given $U\subset V$, we will use $\phi_{U}(x_{I(U)})$ to denote the configuration $\left(\phi_{u}\left(x_{I(u)}\right)\right)_{u\in U}\in\mathtt{B}^{U}$.
@@
The couple $(\mathtt{B}^{V},F)$ is a finite dynamical system whose behavior can be encompassed in a directed graph with vertices in $\mathtt{B}^{V}$ and arrows $x\mapsto y$ for each couple $(x,y)\in\mathtt{B}^{V}\times\mathtt{B}^{V}$ such that $F(x)=y$. This directed graph $\mathcal{T}_{F}$ is the transition diagram of the dynamical system. It is easy to verify that each connected component $C$ of $\mathcal{T}_{F}$ consists of a single cycle decorated with a certain number of directed trees rooted at a point in the cycle. Each cycle codifies a periodic attractor of $(\mathtt{B}^{V},F)$, and the connected component containing this cycle codifies its entire basin of attraction. In this context, the attractors’ landscape of $(\mathtt{B}^{V},F)$ is nothing but the structure of its transition diagram $\mathcal{T}_{F}$. We will characterize this structure by a list containing, for each connected component $C$, the length of the cycle $P(C)$ (i.e., the period of the attractor), the size of the component $|C|$ (i.e., the size of the basin of attraction), the mean length $\langle\tau\rangle_{C}$ and the maximal length $\tau_{\max}(C)$ of the attached trees.
@@
Consider a network $(V,A)$ such that $I(v)\neq\emptyset$ for each $v\in V$. We say that $U\subset V$ determines the set $U^{\prime}\subset V$, if $I(U^{\prime})\subset U$, i.e., if the input set of each node in $U^{\prime}$ is completely contained in $U$. Given $U\subset V$, with $\partial U$ we denote the maximal set (in the sense of containment) determined by $U$, i.e., $\partial U=U^{\prime}\Leftrightarrow(I(U^{\prime})\subset U\,\wedge\,I(W)\subset U\Rightarrow W\subset U)$. The set $U\subset V$ is dominant if the chain $U_{0}:=U\subset U_{1}\subset\cdots\subset U_{d}:=V$, where $U_{\ell+1}=U_{\ell}\cup\partial U_{\ell}$ for each $0\leq\ell\leq d$. We call $U_{0}\subsetneq\cdots\subsetneq U_{d}$ the chain determined by $U$. Given a dominant set $U$, its chain is uniquely determined and we refer to its length $d=d(U)$, as the depth of the dominant set $U$. Alternatively, we can characterize the dominant sets is as follows.
@@
Figure 1. A five-vertex network with a dominant set of two vertices colored in blue, and depth $d=2$. The reduced graph is shown aside.
Remark on dominant sets
The definition of dominant sets of vertices depends only on the topology of the underlying network and makes sense for Boolean networks with asynchronous updating. A given network admits multiple dominant sets: already, supersets of a dominant set are dominant as well. Even restricted to sets of minimal size, dominant sets are not unique. This can be easily verified by considering the directed cyclic graph, in which any one-vertex subset is dominant and of minimal size. Although dominant sets are easily characterized, identifying minimal dominant sets is, in general, computationally hard. In our contexts, finding a dominant set of minimal size is equivalent to finding a minimal feedback vertex set in a directed graph, a problem known to be NP-complete [17].
@@ 3. The induced system
    Since the dominant set is not unique, the induced network could in principle depend on the choice of the dominant set. Natural choices of dominant sets are dominant sets of minimal cardinality, but even in that case, two dominant sets with the same minimal cardinality could have different recurrence lengths and non-equivalent reduced networks.
@@
    The eventual conjugacy is an equivalence relation between dynamical systems on finite sets. Reflexivity and transitivity are obvious. To prove symmetry, consider the following:
@@
Consider the Boolean network $(\mathtt{B}^{V},F)$ with underlying graph $(V,A)$. Let $U\subset V$ be a dominant set with depth $d=d(U)$, and recurrence length $\ell=\ell(U)$. Let $\left(\mathtt{B}_{\ell}^{U},\mathcal{F}\right)$ be the induced automata network, and $\mathtt{h}:\mathtt{B}^{V}\to\mathtt{B}_{\ell}^{U}$ the eventual conjugacy between both systems. Recall that ${\rm Per}_{P}(F)$ denotes the set of all $F$-periodic points of the period $P$ and ${\rm Per}(F)=\bigcup_{P\in\mathbb{N}}{\rm Per}(F)$. For each $x\in(\mathtt{B}^{V},F)$, with $\tau_{F}(x)$ we denote the transient time of the orbit starting at $x$, that is,
@@
As mentioned above, the landscape of the attractors comprises a finite number of basins attached to the periodic attractors, each corresponding to a connected component of the transition diagram $\mathcal{T}_{F}$. For each $x\in\mathtt{B}^{V}$, denote by $C_{F}(x)$ the basin of attraction containing $x$. Finally, let us denote by $N_{\mathcal{A}}$ the number of attractors and by $N_{\mathcal{A}}(P)$ the number of those attractors of minimal period $P$. The same notions and corresponding notations apply to the induced automata network $(\mathtt{B}_{\ell}^{U},\mathcal{F})$. We have the following results.
@@ 4. Clover networks with signed majority rule
A local rule of this kind is a signed majority rule, with signs given by the arrow map $\sigma$. In this case, to each directed cycle $C=v_{0}\mapsto v_{1}\mapsto\cdots\mapsto v_{m}\mapsto v_{0}$ in $(V,A)$, we associate the sign
@@
Let $(V,A)$ be a clover network and $\sigma:A\to\mathtt{B}$ the interaction signs. Let $\mathcal{S}:=(\sum_{|C|=t}\sigma(C):1\leq t\leq\ell)$ be the vector formed by the cycles’ sign sums, grouped according to their length. Then, the corresponding Boolean network with signed majority rule induces an automata network such that $\mathcal{F}\left(y^{0}y^{1}\cdots y^{\ell-1}\right)=\left(\Phi(y)y^{0}y^{1}\cdots y^{\ell-1}\right)$, with $\Phi:\mathtt{B}_{\ell}\to\mathtt{B}$ such that
@@
The recursion stops at $t=\ell-1$, with $\ell=\max\{|C|:\,C\text{ is a cycle}\}$. The clover structure ensures that each $u\in I(v_{0})$ appears only in one cycle $C_{u}=v_{0}\mapsto\cdots\mapsto u\mapsto v_{0}$, and for that cycle, the recursion above gives $x_{u}^{0}=\sigma(C_{u})y^{|C_{u}|-1}$, from which the proposition follows. $\Box$
@@
    The dynamics $F:\mathtt{B}^{V}\to\mathtt{B}^{V}$, defined by Equation (7), commutes with the symmetry $x\mapsto-x$ in $\mathtt{B}^{V}$. This derives from the fact that ${\rm Sign}(-x|\,-y)=-{\rm Sign}(x,y)$ for each $x\in\mathbb{Z},\,y\in\mathtt{B}$. The induced automata, $\mathcal{F}:\mathtt{B}_{\ell}\to\mathtt{B}_{\ell}$ inherits this symmetry.
@@
    The symmetry $x\mapsto-x$ reflects in a symmetry in the transition diagram for both, the original and the induced automatq.
@@
    Suppose that $\mathcal{S}$ is such that $|\mathcal{S}_{\tau}|>\sum_{t\neq\tau}^{\ell}|\mathcal{S}_{\tau}|$ for some $1\leq\tau\leq\ell$. In this case the induced automata network is such that $y\mapsto\left({\rm Sign}(\mathcal{S}_{\tau}\,y^{\tau-1})y^{0}\cdots y^{\ell-2}\right)$ and we have two possibilities:
@@
In what follows we analyze the landscape of attractors in an ensemble of clover Boolean networks with signed majority rule. We compare the mean structure of the attractor’s landscape of the original Boolean network to that of induced automata network. The aim is to evaluate the influence of the sign distribution and the distribution of cycle lengths on key dynamical features such as the number of attractors, their periods, the size of their respective basins, and the transient behavior preceding convergence. At the same time we will study the effect of network reduction, on the transient behavior. In this way we can asses who loose are the theoretical bounds established in Corollary 1, concerning the transient lengths and the size of the basins of attraction.
@@
In this way we obtain a clover network with cycles $1\mapsto n_{k}\mapsto(n_{k}+1)\mapsto\cdots\mapsto(n_{k+1}-1)\mapsto 1$, for $1\leq k\leq s+1$. The parameter $p$ controls the number and the size of the of cycles of the clover network, and we will refer to it as the folding probability. Indeed, the number $N_{c}$ of cycles is distributed according to $\mathbb{P}(N_{c}=k)=1+f(k,N-2,p)$, with f(k,N2,p)f(k,N-2,p) the binomial distribution with N2N-2 trials and parameter $p$. Hence, the expected number of cycles is $\mathbb{E}(N_{c})=1+p(N-2)$. The length $\ell_{1}$ of the cycle $C_{1}$ is such that $\mathbb{P}(\ell_{1}=\ell)=p^{\chi_{<N}(\ell)}\,(1-p)^{\ell-2}$, with $\chi_{<N}$ is the characteristic function of the set $\{\ell<N\}$. This gives an expected length $\mathbb{E}(\ell_{1})=\left(-p^{2}-(1-p)^{N}\right)/\left(p(1-p)\right)\rightarrow(1+p)/p$, as $N\to\infty$. This asymptotically coincides with the common length of the cycles in a clover network with 1+p(N2)1+p(N-2) cycles of the same length. We can estimate the distribution of the maximal length $\max_{1\leq t\leq\ell}\ell_{t}$ by considering that our random structure is reasonably modeled by a clover network with $N_{c}=1+p(N-2)$ cycles whose lengths are i.i.d. random variables with exponential distribution $\mathbb{P}(\ell_{k}=\ell)=p(1-p)^{\ell-2}$. For this simplified model,
@@
    According to Equation (11), the influence of the inhibition probability becomes less important as the size of the cycles grows. Hence, the large coordinates of $\mathcal{S}$ can be modeled by centered random variables with a variance tending to zero.
@@
    The common number $N_{\mathcal{A}}$ of connected components of $\mathcal{T}_{F}$ and $\mathcal{T}_{\mathcal{F}}$, i.e., the common number of attractors.
@@
    The common length $P(\mathcal{C})$ of the cyclic core of each component of $\mathcal{T}_{F}$ and $\mathcal{T}_{\mathcal{F}}$, i.e., the common period of the corresponding periodic attractors.
@@
    The reduction after induction of the mean depth $\Delta\tau=\langle\tau_{F}\rangle_{C_{F}}-\langle\tau_{\mathcal{F}}\rangle_{\mathcal{C}_{\mathcal{F}}}$ of each component, which is noting but the difference between the basins’ mean transients before and after induction.
@@
    The reduction after induction of the maximal depth $\Delta\tau_{\max}=\max_{x\in C}\tau_{F}(x)-\max_{y\in\mathtt{h}(C)}\tau_{\mathcal{F}}(y)$ of each component, which is noting but the difference between the basins’ maximal transients before and after induction.
We register as well, the recurrence length $\ell$, of the underlying clover network, which serves a key parameter for the theoretical upper bounds established in Corollary 1.
We performed 500 simulations with $N=10$, for each $p,q\in\{0.3,0.6,0.9\}$, observing some interesting features regarding the attractors’ landscape. Notice that the total number of admissible clover configurations is bounded above by $2\cdot 3^{N-2}=2\cdot 3^{8}$. In the case of normally distributed observables, this number of realizations ensures, in the worst-case scenario, a confidence level of $95\%$ with a margin of error below $5\%$ for proportion-type observables.
@@
    The number $N_{\mathcal{A}}$ of attractors was, in all of our simulations, much smaller than the theoretical upper bound $\sum_{P=1}^{P_{\max}}2^{P}/P$, which indicates that the number of attractors of a given period, $N_{\mathcal{A}}(P)$, is usually smaller than the theoretical upper bound $2^{P}/P$. Besides, $N_{\mathcal{A}}$ seems to decrease with the folding probability $p$, consistent with the fact that $N_{\mathcal{A}}$ increases with the recurrence length $\ell$. For small $p$, the distribution of $N_{\mathcal{A}}$ is highly skewed and exhibits a wide $q_{0.05},q_{0.95}]$ interval, which captures the central 90% of the realizations and therefore reflects substantial dispersion. For larger $p$ the interval contracts significantly, indicating a clear concentration phenomenon.
@@
    0.3 26.5  [1,52] 3.0  [1,5] 2.0  [1,3]
    0.6 22.5  [1,44] 3.5  [1,6] 2.0  [1,3]
    0.9 19.0  [2,36] 3.5  [2,5] 3.0  [3,3]
    $\ell$ 5.87 $\pm$ 1.71 4.06 $\pm$ 1.07 2.66 $\pm$ 0.67
    Table 1. The median number of attractors, $N_{\mathcal{A}}$, together with the empirical $q_{0.05},q_{0.95}]$ interval over 500 simulations, as a function of the folding and inhibition probabilities. In the last line, the recurrence length $\ell$ is reported as mean $\pm$ standard deviation. All simulations were performed with $N=10$ nodes.